OpenStudy (anonymous):

A bag has 4 yellow, 6 red, 2 green, and 8 purple marbles. What is the probability of picking a purple marble, replacing it, and then picking another purple marble?

OpenStudy (anonymous):

Since you are replacing the marble after the first pick, the events are independent of each other and your total probability is the probability of picking a purple marble, squared. P(first purple) is: (# of purple marbles)/(total # of marbles) Now just square that and you have your answer.

OpenStudy (anonymous):

0.16?

OpenStudy (anonymous):

4/25

OpenStudy (anonymous):

P(first purple) is: (# of purple marbles)/(total # of marbles) = 8/20 = 2/5 (2/5)^2 = 4/25 Yes, you got it! Good job!

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition! @lxygrl

OpenStudy (anonymous):

And, as you know, that's the same as 0.16. Good job!